OpenStudy (anonymous):

A cyclist takes (x+3) hours to travel a distance of 108km when cycling at x km per hour. Find: a) the value of x b) the time taken to complete the journey

OpenStudy (dls):

x=108/x+3 (s=d/t)

OpenStudy (dls):

solve for x,you will get a quadratic equation in x

OpenStudy (anonymous):

x=-9 or x=12? and what about part b?

OpenStudy (dls):

so x=12=speed you just wrote time=x+3 time=12+3=15hours

OpenStudy (anonymous):

negative value of x must be rejected